# Bluesky: Network Topology, Polarization, and Algorithmic Curation

### v2

@@ Abstract
Keywords Decentralized Social Media  $\cdot$ Network Topology  $\cdot$ Temporal Network Analysis  $\cdot$ Polarization  $\cdot$ Personalized Content Curation
@@ Results
Fig 1 presents the daily number of active users according to six engagement metrics. Dates with the highest numbers of new users over the year 2023 were substantially driven by activity and news about X (formerly Twitter). While further research leveraging qualitative surveys is necessary to establish the exact reasons for users switching, the number of sign-ups significantly correlates with news about Twitter [31].
@@
Table 1: User activity distributions are indicative of heavy-tailed behavior. Included metrics are the mean $\mu$, standard deviation $\sigma$, skewness $\gamma$, kurtosis $\beta$, minimum $m$, maximum $M$, and the ratio of mean to maximum $\frac{\mu}{M}$ from the distributions depicted in Fig 2 and 3. The exponent of the power law distribution is denoted by $\alpha$ The notably high values of $\gamma$ and $\beta$ indicate a pronounced right-skewed, heavy-tailed nature across all distributions. Furthermore, the exceptionally low $\frac{\mu}{M}$ values further confirm the extensive tail behavior characteristic of these distributions. ”#” should be read as Number of.
@@
[Uncaptioned image] Adult Content Formerly “Suggestive” All (nonviolent) … 4,180
@@
[Uncaptioned image] New Posts by furries across Bluesky… 3,221
@@
[Uncaptioned image] Art Posts by furries with #furryart… 2,938
@@
The most liked feed For You has been renamed Discover and promises to show ”Trending content from your personal network” [5]. Other popular feeds include ”Science”, which is a feed curated by ”Bluesky professional scientists” [4]. Other feeds such as the ”Hospitality & Tourism [Uncaptioned image]” or ”Paleo Sky” use regex patterns to match posts (tourism, skift and Paleontology|Archaeology|#PaleoSky respectively).
@@
Table 3: Activity related to Feeds on Bluesky. Included metrics are the mean $\mu$, standard deviation $\sigma$, skewness $\gamma$, kurtosis $\beta$, $m$, maximum $M$, and the ratio of mean to maximum $\frac{\mu}{M}$ from the distributions depicted in Fig 2 and 3. $\alpha$ indicates the exponent of the power law distribution. The notably high values of $\gamma$ and $\beta$ indicate a pronounced right-skewed, heavy-tailed nature across all distributions. Furthermore, the exceptionally low $\frac{\mu}{M}$ values further confirm the extensive tail behaviour characteristic of these distributions.
@@ Discussion
First, we extracted all posts containing the keywords “Israel”, “Palestine” and their translations into languages present in the dataset (Arabic, English, French, German, Greek, Hebrew, Italian, Japanese, Korean, Persian, Russian, Ukrainian, Azerbaijani, Danish, Dutch, French, Finnish, Hungarian, Indonesian, Kazakh, Chinese, Norwegian, Portuguese, Romanian, Slovene, Spanish, Swedish, Tajik, Turkish). Subsequently, we calculated the mutual information of all uni-, bi-, and tri-grams from the dataset with any of initial seed terms, filtering for a minimal number of occurrences of above 50 times. We then manually reviewed the top 100 n-grams, ranked by their mutual information, extracted for each of the initial seed terms and selected all n-grams directly related to the conflict. Lastly, to prevent any biases across languages, we translated all selected n-grams into all the initial languages to ensure that no variation in n-grams included bias the distribution of stance scores. We then queried the database for any of the retrieved n-grams and created a dataset of 1.3 Million posts related to the conflict. From this subset, we manually annotated a random sample of 1,000 posts. Each post was labeled as Pro-Israel (1), Neutral (0), or Pro-Palestine (-1) based on the stance expressed in the content. An additional set of 1.000 posts was labeled by crowdworkers via Appen.com. Crowdworkers were presented with posts from Bluesky and prompted with the following question. “Determine the stance (favor Israel, favor Palestine, or neither) of the author towards the Israel/Palestine conflict from each given social media post.” Each participant was given examples for each of the categories and additionally tested on ten quality assurance questions. The average agreement between annotators was 71.23%. We additionally removed any posts where no majority opinion among annotators emerged.
@@
Stance prediction involves automatically determining the position or attitude expressed in a piece of text toward a specific target or topic. We use the multilingual transformer-based language modelXLM-RoBERTa large [10], which is well-suited for the stance prediction task due to its ability to capture cross-lingual semantic information. The data is preprocessed by tokenizing the text using the XLM-RoBERTa tokenizer, with a maximum sequence length of 128 tokens. The model is trained on an A100 GPU and evaluated on a hold-out test set. The final training configuration is reached via fine-tuning had a dropout-probability of 0.14, a learning rate of 2.2$\times$ 10-05 and a weight decay of 0.001. In addition, we freeze the first 21 (out of 24) layers of the model to reduce the chance of overfitting and accelerate training. The best model is finally determined by maximizing the Macro F1 score. Predictions are obtained by applying the trained model to the tokenized test data and selecting the class with the highest probability. Classification metrics, including precision, recall, and F1-score, are reported for each class in table 6. The model significantly outperforms random guessing and a simple benchmarking naive bayes model which achieves a macro F1 score of 0.47 shown in table 7.
@@
Ethics Statement
This study was approved by the Ethics Committee of the Faculty of Philosophy at the University of Zurich for the period of March 6, 2023, to February 28, 2027. As the research involved only the collection of publicly available retrospective data from online platforms, informed consent was not obtained from individual participants. Only openly accessible data was collected. The study was conducted in the online public space, and no private or protected data were accessed. To protect participant privacy, raw data will not be publicly released.

### v3

@@ Results
Figure 1 presents the daily number of active users according to six engagement metrics. Dates with the highest numbers of new users over the year 2023 were substantially driven by activity and news about X (formerly Twitter). While further research leveraging qualitative surveys is necessary to establish the exact reasons for users switching, the number of sign-ups significantly correlates with news about Twitter [17].
@@
Table 1: User activity distributions are indicative of heavy-tailed behavior. Included metrics are the mean $\mu$, standard deviation $\sigma$, skewness $\gamma$, kurtosis $\beta$, minimum $m$, maximum $M$, and the ratio of mean to maximum $\frac{\mu}{M}$ from the distributions depicted in Fig. 2 and 3. The exponent of the power law distribution is denoted by $\alpha$ The notably high values of $\gamma$ and $\beta$ indicate a pronounced right-skewed, heavy-tailed nature across all distributions. Furthermore, the exceptionally low $\frac{\mu}{M}$ values further confirm the extensive tail behavior characteristic of these distributions. ”#” should be read as Number of.
@@
Adult Content Formerly “Suggestive” All (nonviolent) … 4,180
@@
New Posts by furries across Bluesky… 3,221
@@
Art Posts by furries with #furryart… 2,938
@@
The most liked feed For You has been renamed Discover and promises to show ”Trending content from your personal network” [9]. Other popular feeds include ”Science”, which is a feed curated by ”Bluesky professional scientists” [29]. Other feeds such as the ”Hospitality & Tourism” or ”Paleo Sky” use regex patterns to match posts (tourism, skift and Paleontology|Archaeology|#PaleoSky respectively).
@@
Table 3: Activity related to Feeds on Bluesky. Included metrics are the mean $\mu$, standard deviation $\sigma$, skewness $\gamma$, kurtosis $\beta$, $m$, maximum $M$, and the ratio of mean to maximum $\frac{\mu}{M}$ from the distributions depicted in Fig. 2 and 3. $\alpha$ indicates the exponent of the power law distribution. The notably high values of $\gamma$ and $\beta$ indicate a pronounced right-skewed, heavy-tailed nature across all distributions. Furthermore, the exceptionally low $\frac{\mu}{M}$ values further confirm the extensive tail behaviour characteristic of these distributions.
@@ Discussion
First, we extracted all posts containing the keywords “Israel”, “Palestine” and their translations into languages present in the dataset (Arabic, English, French, German, Greek, Hebrew, Italian, Japanese, Korean, Persian, Russian, Ukrainian, Azerbaijani, Danish, Dutch, French, Finnish, Hungarian, Indonesian, Kazakh, Chinese, Norwegian, Portuguese, Romanian, Slovene, Spanish, Swedish, Tajik, Turkish). Subsequently, we calculated the mutual information of all uni-, bi-, and tri-grams from the dataset with any of initial seed terms, filtering for a minimal number of occurrences of above 50 times. We then manually reviewed the top 100 n-grams, ranked by their mutual information, extracted for each of the initial seed terms and selected all n-grams directly related to the conflict. Lastly, to prevent any biases across languages, we translated all selected n-grams into all the initial languages to ensure that no variation in n-grams included bias the distribution of stance scores. A full list of all terms with at least 1.000 exclusive posts (posts not added by any other n-grams), can be found in table 8 in the appendix. We then queried the database for any of the retrieved n-grams and created a dataset of 1.3 Million posts related to the conflict. From this subset, we manually annotated a random sample of 1,000 posts. Each post was labeled as Pro-Israel (1), Neutral (0), or Pro-Palestine (-1) based on the stance expressed in the content. An additional set of 1.000 posts was labeled by crowdworkers via Appen.com. Crowdworkers were presented with posts from Bluesky and prompted with the following question. “Determine the stance (favor Israel, favor Palestine, or neither) of the author towards the Israel/Palestine conflict from each given social media post.” Each participant was given examples for each of the categories and additionally tested on ten quality assurance questions. The average agreement between annotators was 71.23%. We additionally removed any posts where no majority opinion among annotators emerged.
@@
Stance prediction involves automatically determining the position or attitude expressed in a piece of text toward a specific target or topic. We use the multilingual transformer-based language model XLM-RoBERTa large [49], which is well-suited for the stance prediction task due to its ability to capture cross-lingual semantic information. The data is preprocessed by tokenizing the text using the XLM-RoBERTa tokenizer, with a maximum sequence length of 128 tokens. The model is trained on an A100 GPU and evaluated on a hold-out test set. The final training configuration is reached via fine-tuning had a dropout-probability of 0.14, a learning rate of 2.2$\times$ 10-05 and a weight decay of 0.001. In addition, we freeze the first 21 (out of 24) layers of the model to reduce the chance of overfitting and accelerate training. The best model is finally determined by maximizing the Macro F1 score. Predictions are obtained by applying the trained model to the tokenized test data and selecting the class with the highest probability. Classification metrics, including precision, recall, and F1-score, are reported for each class in table 6. The model significantly outperforms random guessing and a simple benchmarking naive Bayes model, which achieves a macro F1 score of 0.47, shown in table 7.
@@
S1 Appendix
Term Translation Total Posts Exclusive Posts
israel 632,737 274,679
gaza 336,061 126,116
hamas 233,000 82,169
palestine 127,002 62,032
jewish 92,716 55,697
antisemitism 59,499 35,680
jews 71,147 34,178
rafa 45,856 29,255
Israel* Israel 51,748 28,096
zionis Zionism 58,451 27,168
palestinians 90,392 26,539
netanyahu 49,334 17,871
judar Jews 14,115 13,575
Israel Israel 15,577 13,277
juden Jews 16,453 9,252
Israel Israel 11,696 9,097
israël Israel 22,843 8,175
Palestine* Palestine 23,045 7,777
jüdisch Jewish 12,058 7,324
sionism Zionism 6,364 5,583
judaism 9,554 4,616
palestina Palestine 8,878 4,435
ebrei Jews 4,256 4,047
Hamas* Hamas 12,883 3,190
Palestine Palestine 5,845 3,149
Hamas Hamas 4,781 3,034
west bank 17,698 2,870
october 7 11,418 2,808
juif Jewish 3,834 2,624
antisémitisme Antisemitism 3,360 2,545
Israel§ Israel 3,172 2,347
palästina Palestine 6,769 2,300
Israel Israel 3,007 2,251
palästinenser Palestinians 7,995 2,004
Hamas Hamas 3,022 1,960
israélien Israeli 6,276 1,952
from the river to the sea 5,485 1,894
Netanyahu* Netanyahu 4,581 1,593
Gaza Strip* Gaza Strip 4,742 1,427
sionisti Zionists 1,376 1,241
gazze Gaza 1,643 1,207
Israel Israel 1,765 1,126
joden Jews 1,599 1,021
Table 8: Selected query terms used to identify posts related to the Israel-Palestine conflict on Bluesky. Terms with at least 1,000 exclusive posts (posts not matched by any other n-gram), derived from an initial set of seed terms (“Israel,” “Palestine,” and their translations) through mutual information calculation and manual review. Terms are listed in descending order of exclusive posts.
* Transliterated from Arabic.
† Transliterated from Hebrew.
‡ Transliterated from Japanese.
§ Transliterated from Korean.
¶ Transliterated from Russian.